# Deep Learning-Based CSI Feedback for Wi-Fi Systems With Temporal Correlation

### v1

@@ Abstract
To achieve higher throughput in next-generation Wi-Fi systems, a station (STA) needs to efficiently compress channel state information (CSI) and feed it back to an access point (AP). In this paper, we propose a novel deep learning (DL)-based CSI feedback framework tailored for next-generation Wi-Fi systems. Our framework incorporates a pair of encoder and decoder neural networks to compress and reconstruct the angle parameters of the CSI. To enable an efficient finite-bit representation of the encoder output, we introduce a trainable vector quantization module, which is integrated after the encoder network and jointly trained with both the encoder and decoder networks in an end-to-end manner. Additionally, we further enhance our framework by leveraging the temporal correlation of the angle parameters. Specifically, we propose an angle-difference feedback strategy which transmits the difference between the current and previous angle parameters when the difference is sufficiently small. This strategy accounts for the periodicity of the angle parameters through proper preprocessing and mitigates error propagation effects using novel feedback methods. We also introduce a DL-based CSI refinement module for the AP, which improves the reconstruction accuracy of the angle parameters by simultaneously utilizing both the previous and current feedback information. Simulation results demonstrate that our framework outperforms the standard method employed in current Wi-Fi systems. Our results also demonstrate significant performance gains achieved by the angle-difference feedback strategy and the CSI refinement module.
@@ I Introduction
Recently, wireless networks have dramatically advanced to meet the growing demand for higher data rates. Specifically, the IEEE 802.11 task group has been developing a novel extremely high throughput (EHT) amendment for wireless local area networks (WLANs) [1]. With developing an advanced multiple-input multiple-output (MIMO) techniques, it aims to service much higher throughput and reliable communications. To achieve higher throughput in next-generation Wi-Fi systems utilizing MIMO techniques, acquiring an accurate downlink channel state information (CSI) is indispensable for the access point (AP) to transmit data to the station (STA). The typical CSI acquisition process in the system is initiated by the AP’s transmitting the null data packet announcement (NDPA) frame to STA. NDPA serves the role of notifying the STA of all necessary probing details of MIMO transmission. Then, a null data packet (NDP) is transmitted to the STA for estimating the downlink CSI. After the channel estimation, the STA compresses and feeds the CSI back to the AP, in the form of the compressed beamforming report (CBR). The compression is performed by extracting angle parameters after performing singular value decomposition (SVD) of the MIMO channel. The aforementioned protocol in the Wi-Fi systems is summarized in Fig. 1. Another approach to acquire the downlink CSI is to leverage channel reciprocity. In this process, the downlink CSI is inferred from the uplink CSI with additional calibrations. However, this approach is excluded from IEEE 802.11ac standard [2], indicating its verified limitation.
@@
The CSI feedback methodologies have been an important problem in cellular networks as well. Among various methodologies, deep learning (DL)-based CSI feedback methods have been proposed for the cellular networks in recent years [3, 4, 5, 6, 7, 8]. By properly preprocessing the CSI and carefully designing neural networks, these methods have demonstrated their effectiveness in reducing CSI feedback overhead, outperforming traditional non-DL-based approaches like compressed sensing and static codebook-based feedback[9, 10]. A typical approach of the DL-based CSI feedback is to compress a preprocessed CSI into a low-dimensional latent vector by using an encoder network. Then, the CSI is reconstructed by using a decoder network from a received latent vector. To enable the transmission of the latent vector with a finite number of bits, one possible approach is to apply scalar quantization (SQ) to each latent element. For example, non-uniform SQ methods were explored in [11, 12] to represent the latent vector with a limited number of bits for CSI feedback in cellular networks. However, the application of SQ to DL-based CSI feedback has inherent limitations. First, SQ disregards the correlations between latent elements, resulting in suboptimal performance when these elements are correlated [13]. Second, SQ cannot allocate fewer than one quantization bit per latent entry, which imposes constraints on the dimensionality of the latent vector for a given feedback overhead and ultimately limits the feedback efficiency. These limitations can be resolved by applying a vector quantization (VQ) in the latent space. The application of VQ to the latent representation allows for modeling the correlation among latent entries. Furthermore, the joint optimization of the VQ codebook with the entire model facilitates quantization that effectively reflects the distribution of the latent vectors, thereby enhancing the efficiency of the learned representations [14]. Motivated by this, DL-based CSI feedback methods incorporating VQ have been proposed for cellular networks in [15, 16, 17]. In [15], it was reported that the DL-based CSI feedback with the VQ framework can achieve a better CSI reconstruction performance than that of DL-based CSI feedback with the SQ framework in cellular networks.
@@
To further improve the efficiency of DL-based CSI feedback, some studies have leveraged the temporal correlation of CSI, which naturally arises in practical communication scenarios involving devices with low to moderate mobility. A common approach in this direction is to employ neural network architectures specifically designed for time-series inputs. For instance, it was reported in [18, 19] that incorporating long short-term memory (LSTM) [20] enables the model to capture temporal correlation, enhancing feedback performance. On the other hand, an alternative approach to leveraging the temporal correlation of CSI is to exploit the inherent sparsity in the differences between sequential CSI instances. This approach stems from the intuition that sequential instances are likely to exhibit similar values, resulting in sparse differences which is more efficient form to compress and reconstruct. The sparsity in CSI differences was leveraged in [22, 21], demonstrating an efficacy of this approach in improving the performance of the CSI feedback in cellular networks.
@@
Unlike DL-based CSI feedback in cellular networks, limited attention has been given to DL-based CSI feedback for Wi-Fi systems [23, 24]. In [23], the angle parameters of downlink CSI were separately compressed using hyperparameterized preprocessing and two distinct fully connected networks (FCNs). In [24], a convolutional neural network (CNN)-based model with a channel attention module was used to compress a right singular matrix in SVD, rather than the angle parameters. However, these studies employ simple quantization techniques without jointly training the quantization module with the encoder and decoder networks. Moreover, none of the existing research has leveraged the temporal correlation of CSIs in DL-based CSI feedback specifically for Wi-Fi systems. The potential benefits of the joint training of the quantization module and the utilization of temporal correlation in CSIs remain unexplored, despite their critical importance in realizing the full potential of DL-based CSI feedback in next-generation Wi-Fi systems.
@@
  • We introduce an initial feedback strategy for our framework, which integrates a trainable VQ module with the DL-based CSI feedback process in Wi-Fi systems. Our strategy basically facilitates joint training of the VQ module with encoder and decoder networks. This joint training enables an efficient finite-bit representation of the latent vector, improving feedback efficiency. Additionally, we determine an appropriate input form for the encoder network by aggregating angle parameters specified in the IEEE 802.11 standard.
  • We propose a novel angle-difference feedback strategy for Wi-Fi systems that leverages the temporal correlation of CSIs, an aspect that has never been explored in DL-based CSI feedback for Wi-Fi systems. The core idea of our framework is to feed back the difference between the current and previous angle parameters when the difference is sufficiently small. To ensure an appropriate input for the angle-difference encoder network, we introduce a preprocessing step designed to capture the periodicity of angle parameters in Wi-Fi systems. Moreover, to address the challenge of error propagation inherent in the angle-difference feedback, we design two VQ architectures, referred to as parallel VQ and unified VQ, each tailored to compensate for residual errors from the previous feedback.
  • We propose a novel CSI refinement framework for Wi-Fi systems that enhances reconstruction performance at the AP by leveraging the temporal correlation of CSIs. A key ingredient of this framework is a CSI refinement module designed to refine the angle parameters at each time instance using both the previous and current feedback information. To further improve its performance, we also introduce a recursive refinement strategy, where the refined angle parameter is utilized as an input for subsequent CSI refinement processes. These advancements provide a new approach to exploiting temporal correlations in CSIs on the AP side as well, further improving the efficiency and accuracy of CSI reconstruction in Wi-Fi systems.
  • Using extensive simulations, we validate the effectiveness of the proposed CSI feedback and refinement frameworks in Wi-Fi systems with temporally-correlated channels. Our evaluations leverage a CSI dataset generated from the IEEE 802.11be standard configuration [1]. The simulation results reveal that the proposed CSI feedback, angle-difference feedback, and CSI refinement frameworks consistently achieve significant performance gains across diverse scenarios. Furthermore, we demonstrate that the proposed frameworks outperforms the standard feedback method employed in current Wi-Fi systems, both in terms of CSI reconstruction accuracy and net throughput performance. Through these results, we demonstrate the potential of our frameworks in advancing the efficiency of next-generation Wi-Fi systems.
@@ II-A Downlink MIMO Wi-Fi Systems
Now, we present the CSI compression method employed in the IEEE 802.11 standard, which is based on the Givens rotation with angle parameters $(\phi,\psi)$[27, 28, 29]. First, the phase offset of the last row of $\overline{\mathbf{V}}[k]$ is eliminated as
@@ 2. Extraction:
These parameters are fed back to the AP in order of $[...,\ \hat{\phi}^{(k)}_{i:N_{t}-1,i},\ \hat{\psi}^{(k)}_{i+1:N_{t},i},\ ...]^{\rm T}$ for the $k$-th subcarrier. The aforementioned feedback method provides satisfactory performance when the number of feedback bits is sufficient, which is why it has been adopted in the IEEE 802.11 standard to date. However, when feedback overhead is limited, this method suffers from significant performance degradation, potentially creating a bottleneck in next-generation Wi-Fi systems. This limitation naturally motivates the development of a new CSI feedback method based on DL.
@@ III-A Initial Feedback Method: DL-based CSI Feedback with VQ
Building on the effectiveness of VQ in facilitating finite-bit representations of latent vectors, we adopt it as the foundation of the inital CSI feedback framework. Our framework utilizes a trainable quantization codebook placed in the latent space and jointly trains the encoder, codebook, and decoder using a loss function that captures both the quantization and reconstruction errors. To avoid a significant computational complexity required for VQ, we adopt the product VQ structure in [26]. Specifically, we first divide the latent vector ${\bf z}$ into $N$ sub-vectors each with dimension $D$ ($M=N\times D$) and then use a $D$-dimensional codebook for quantizing each sub-vector separately, as illustrated in Fig. 2. Let $\mathcal{B}$ be a vector codebook using $B$ bits which consists of $2^{B}$ $D$-dimensional codewords, namely $\{\mathbf{b}_{k}\}^{2^{B}}_{k=1}$. Also, let ${\bf z}_{i}$ be the $i$-th sub-vector of ${\bf z}$, defined as ${\bf z}_{i}=[z_{(i-1)D+1},\cdots,z_{iD}]$, where $z_{j}$ is the $j$-th entry of ${\bf z}$. Then each sub-vector ${\bf z}_{i}$ is quantized to $\mathbf{z}_{q,i}$ using the codebook $\mathcal{B}$ according to the minimum Euclidean distance criterion, i.e.,
@@
In our method, we choose an angle parameter $\mathbf{\Phi}\in\mathbb{R}^{2\times N_{a}\times N_{c}}$ as the input $\mathbf{X}$ of the encoder network, which consists of the angle parameters $\{\phi_{i,j}\}$ and $\{\psi_{l,i}\}$, where $N_{a}={\rm max}(\sum^{N_{t}-1}_{i={\rm max}(N_{t}-N_{s},1)}i,1)$ represents the number of each angle parameter per subcarrier. We aggregate the angle parameters in $\mathbf{\Phi}$ in the same order as specified in the IEEE 802.11 standard. The rationale behind this input choice is that we can exploit the abstracted information in the beamforming matrix with orthogonal columns, while effectively accounting for the correlations between these elements. Otherwise, if we were to use ${\bf V}$ as the input instead, the input dimension would typically be larger, and an additional enforcement process would be required to ensure the output has orthogonal columns explicitly. This rationale is numerically justified in Sec. V-B.
@@
Figure 3: An illustration of preprocessing on $\mathbf{\Phi}^{\rm diff}_{t+1}$ designed to capture the periodicity of angle parameters.

### v2

@@ Abstract
To support higher throughput in next-generation Wi-Fi systems, efficient compression and feedback of channel state information (CSI) from a station (STA) to an access point (AP) is essential. This paper proposes a deep learning (DL)-based CSI feedback framework tailored for Wi-Fi systems. The framework employs encoder and decoder networks to compress and reconstruct CSI angle parameters, with a trainable vector quantization (VQ) module enabling efficient finite-bit representation through end-to-end training. To further enhance performance, we introduce an angle-difference feedback strategy that exploits the temporal correlation of the angle parameters by feeding back the difference between the current and previous values. This is complemented by preprocessing that handles the periodicity of angles and tailored VQ modules that compensate for residual quantization errors. Additionally, we present a DL-based CSI refinement module at the AP, which improves reconstruction by jointly using current and prior feedback. Simulation results show that the proposed framework outperforms both standard Wi-Fi feedback and existing DL-based feedback methods, with notable gains from both angle-difference feedback and CSI refinement.
@@ I Introduction
Recently, wireless networks have dramatically advanced to meet the growing demand for higher data rates. Specifically, the IEEE 802.11 task group has been developing a novel extremely high throughput (EHT) amendment for wireless local area networks (WLANs) [1]. With developing an advanced multiple-input multiple-output (MIMO) techniques, it aims to service much higher throughput and reliable communications. To achieve higher throughput in next-generation Wi-Fi systems utilizing MIMO techniques, acquiring an accurate downlink channel state information (CSI) is indispensable for the access point (AP) to transmit data to the station (STA). The typical CSI acquisition process in the system is initiated by the AP’s transmitting the null data packet announcement (NDPA) frame to STA. NDPA serves the role of notifying the STA of all necessary probing details of MIMO transmission. Then, a null data packet (NDP) is transmitted to the STA for estimating the downlink CSI. After the channel estimation, the STA compresses and feeds the CSI back to the AP, in the form of the compressed beamforming report (CBR). The compression is performed by extracting angle parameters after performing singular value decomposition (SVD) of the MIMO channel[2, 3, 1]. The aforementioned protocol in the Wi-Fi systems is summarized in Fig. 1. Another approach to acquire the downlink CSI is to leverage channel reciprocity. In this process, the downlink CSI is inferred from the uplink CSI with additional calibrations. However, this approach is excluded from IEEE 802.11ac standard [2], indicating its verified limitation.
@@
The CSI feedback methodologies have been an important problem for massive MIMO systems in cellular networks as well. Among various methodologies, deep learning (DL)-based CSI feedback methods have been proposed for the cellular networks in recent years [4, 5, 6, 7, 8, 9]. By exploiting the sparsity of CSI in massive MIMO systems and carefully designing neural networks, these methods have demonstrated their effectiveness in reducing CSI feedback overhead, outperforming traditional non-DL-based approaches like compressed sensing and static codebook-based feedback[10, 11]. A typical approach of the DL-based CSI feedback is to compress a preprocessed CSI into a low-dimensional latent vector by using an encoder network. Then, the CSI is reconstructed by using a decoder network from a received latent vector. To enable the transmission of the latent vector with a finite number of bits, one possible approach is to apply scalar quantization (SQ) to each latent element. For example, non-uniform SQ methods were explored in [12, 13] to represent the latent vector with a limited number of bits for CSI feedback in cellular networks. However, the application of SQ to DL-based CSI feedback has inherent limitations. First, SQ disregards the correlations between latent elements, resulting in suboptimal performance when these elements are correlated [14]. Second, SQ cannot allocate fewer than one quantization bit per latent entry, which imposes constraints on the dimensionality of the latent vector for a given feedback overhead and ultimately limits the feedback efficiency. These limitations can be resolved by applying a vector quantization (VQ) in the latent space. The application of VQ to the latent representation allows for modeling the correlation among latent entries. Furthermore, the joint optimization of the VQ codebook with the entire model facilitates quantization that effectively reflects the distribution of the latent vectors, thereby enhancing the efficiency of the learned representations [15]. Motivated by this, DL-based CSI feedback methods incorporating VQ have been proposed for cellular networks in [16, 17, 18]. In [16], it was reported that the DL-based CSI feedback with the VQ framework can achieve a better CSI reconstruction performance than that of DL-based CSI feedback with the SQ framework in cellular networks.
@@
To further improve the efficiency of DL-based CSI feedback in massive MIMO systems, some studies have leveraged the temporal correlation of CSI, which naturally arises in practical communication scenarios involving devices with low to moderate mobility [19, 20, 22, 21, 23, 24, 25]. A common approach in this direction is to employ neural network architectures specifically designed for time-series inputs. For instance, it was reported in [19, 20] that incorporating long short-term memory (LSTM) [26] enables the model to capture temporal correlation, enhancing feedback performance. On the other hand, an alternative approach to leveraging the temporal correlation of CSI is to exploit the inherent sparsity in the differences between sequential CSI instances. This approach stems from the intuition that sequential instances are likely to exhibit similar values, leading to sparse differences that are more efficient to compress and reconstruct. The sparsity in CSI differences was leveraged in [22, 21], demonstrating an efficacy of this approach in improving the performance of the CSI feedback in cellular networks.
@@
Unlike DL-based CSI feedback in cellular networks, limited attention has been given to DL-based CSI feedback for Wi-Fi systems [27, 28]. In fact, existing DL-based CSI feedback methods for cellular systems are developed under massive MIMO assumptions and rely on sparsity of CSI in the angular-delay domain, which does not hold in typical Wi-Fi systems. Furthermore, these methods do not align with current Wi-Fi system protocols, making their direct application much more impractical. From these points, in [27], the angle parameters of downlink CSI in Wi-Fi systems were separately compressed using hyperparameterized preprocessing and two distinct fully connected networks (FCNs). In [28], a convolutional neural network (CNN)-based model with a channel attention module was used to compress a right singular matrix in SVD, rather than the angle parameters. However, these studies employ simple quantization techniques without jointly training the quantization module with the encoder and decoder networks. Moreover, none of the existing research has leveraged the temporal correlation of CSIs in DL-based CSI feedback specifically for Wi-Fi systems. The potential benefits of the joint training of the quantization module and the utilization of temporal correlation in CSIs remain unexplored, despite their critical importance in realizing the full potential of DL-based CSI feedback in next-generation Wi-Fi systems.
@@
  • We present a novel DL-based CSI feedback framework tailored for Wi-Fi systems, which incorporates both initial and angle-difference feedback strategies. For the initial feedback, we determine an appropriate input representation by aggregating angle parameters defined in the IEEE 802.11 standard and employ a trainable VQ module jointly optimized with the encoder and decoder to enable efficient finite-bit compression. For the angle-difference feedback, we exploit the temporal correlation of CSI by feeding back the difference between consecutive angle parameters. To ensure accurate encoding of these differences, we introduce a preprocessing step that accounts for the periodicity of angles, and design two VQ architectures—parallel VQ and unified VQ—to mitigate error accumulation from previous feedback steps.
  • We propose a new CSI refinement framework for Wi-Fi systems that enhances reconstruction performance at the AP by leveraging the temporal correlation of CSIs. We refine angle parameters using both current and previous feedback, and its performance is further improved by a recursive strategy that feeds refined outputs back into future refinements. This framework enables more accurate and efficient CSI reconstruction on the AP side.
  • Using extensive simulations, we validate the effectiveness of the proposed CSI feedback and refinement frameworks for Wi-Fi systems with temporally correlated channels. The results show that the proposed frameworks consistently deliver significant performance gains across various scenarios. Moreover, the proposed frameworks outperform the standard feedback method used in current Wi-Fi systems in terms of both CSI reconstruction accuracy and net throughput.
@@ II-A Downlink MIMO Wi-Fi Systems
Now, we present the CSI compression method employed in the IEEE 802.11 standard, which is based on the Givens rotation with angle parameters $(\phi,\psi)$[32, 33, 34]. Givens rotation is a technique that expresses a matrix with orthonormal columns as a product of successive plane rotations, and it has been utilized not only for CSI parameter extraction but also in a range of applications, including beamforming methodologies[35, 36]. First, the phase offset of the last row of $\overline{\mathbf{V}}[k]$ is eliminated as
@@
Figure 2: An illustration of overall CSI feedback procedures for IEEE 802.11 standard and DL-based methods.
@@ 2. Extraction:
These parameters are fed back to the AP in order of $[...,\ \hat{\phi}^{(k)}_{i:N_{t}-1,i},\ \hat{\psi}^{(k)}_{i+1:N_{t},i},\ ...]^{\rm T}$ for the $k$-th subcarrier. The aforementioned feedback method provides satisfactory performance when the number of feedback bits is sufficient, which is why it has been adopted in the IEEE 802.11 standard to date. However, when feedback overhead is limited, this method suffers from significant performance degradation, potentially creating a bottleneck in next-generation Wi-Fi systems. This limitation naturally motivates the development of a new CSI feedback method based on DL. The aforementioned IEEE standard feedback procedure is illustrated in Fig. 2, alongside a DL-based alternative, which will be discussed in the following sections.
@@ III-A Initial Feedback Method: DL-based CSI Feedback with VQ
Building on the effectiveness of VQ in facilitating finite-bit representations of latent vectors, we adopt it as the foundation of the inital CSI feedback framework. Our framework utilizes a trainable quantization codebook placed in the latent space and jointly trains the encoder, codebook, and decoder using a loss function that captures both the quantization and reconstruction errors. To avoid a significant computational complexity required for VQ, we adopt the product VQ structure in [30, 31]. Specifically, we first divide the latent vector ${\bf z}$ into $N$ sub-vectors each with dimension $D$ ($M=N\times D$) and then use a $D$-dimensional codebook for quantizing each sub-vector separately, as illustrated in Fig. 3. Let $\mathcal{B}$ be a vector codebook using $B$ bits which consists of $2^{B}$ $D$-dimensional codewords, namely $\{\mathbf{b}_{k}\}^{2^{B}}_{k=1}$. Also, let ${\bf z}_{i}$ be the $i$-th sub-vector of ${\bf z}$, defined as ${\bf z}_{i}=[z_{(i-1)D+1},\cdots,z_{iD}]$, where $z_{j}$ is the $j$-th entry of ${\bf z}$. Then each sub-vector ${\bf z}_{i}$ is quantized to $\mathbf{z}_{q,i}$ using the codebook $\mathcal{B}$ according to the minimum Euclidean distance criterion, i.e.,
@@
In our method, we choose an angle parameter $\mathbf{\Phi}\in\mathbb{R}^{2\times N_{a}\times N_{c}}$ as the input $\mathbf{X}$ of the encoder network, which consists of the angle parameters $\{\phi_{i,j}\}$ and $\{\psi_{l,i}\}$, where $N_{a}={\rm max}(\sum^{N_{t}-1}_{i={\rm max}(N_{t}-N_{s},1)}i,1)$ represents the number of each angle parameter per subcarrier. For implementation flexibility, we also allow $\mathbf{\Phi}$ to be reshaped into a flattened format $\mathbf{\Phi}\in\mathbb{R}^{1\times 2N_{a}\times N_{c}}$, where the two components are concatenated along the angle parameter dimension. We aggregate the angle parameters in $\mathbf{\Phi}$ in the same order as specified in the IEEE 802.11 standard. The rationale behind this input choice is that we can exploit the abstracted information in the beamforming matrix with orthogonal columns, while effectively accounting for the correlations between these elements. In contrast, using the full matrix $\mathbf{V}$ as the input typically results in a much higher input dimension and requires additional enforcement to ensure orthogonality in the output. This rationale is numerically justified in Sec. V-B. Ultimately, the proposed DL-based CSI feedback method replaces the uniform quantization used in the conventional standard with an encoder, VQ module, and decoder following the angle parameter extraction process, as illustrated in Fig. 2.
@@
Figure 4: An illustration of preprocessing on $\mathbf{\Phi}^{\rm diff}_{t}$ designed to capture the periodicity of angle parameters.